# Zhou Zhang Zhao Wang Pan Ng Zhang Ni Ma Nguyen Yip Ma Alibaba GroupSingapore RochesterUnited States of America

### v2

@@ 3.4 Subjective Evaluation
We conducted listening experiments, with 10 native English speakers each listening to a total of 260 synthesized speech samples. We first evaluate our proposed framework on the zero-shot TTS task, where the ED is predicted based on the prompt speech (‘Proposed’). We selected ‘CosyVoice’ [13] as the baseline. A mean opinion score (MOS) test is conducted to evaluate the speech naturalness, where participants were asked to score each sample on a 5-point scale (5: Excellent, 4: Good, 3: Satisfactory, 2: Unsatisfactory, 1: Bad). As shown in Table 2, our proposed framework with ED prediction achieves a score of 4.54, outperforming the baseline score of 4.36, while lower than the ground-truth score of 4.80. It shows that our proposed framework with ED prediction achieves better naturalness than the baseline CosyVoice.
@@
We additionally conducted a mean opinion score to assess emotional intelligibility (E-MOS), where participants were asked to rate each speech sample in terms of emotion perception. We evaluated our proposed framework against two baselines: MixedEmotion [25] and CosyVoice [13]. As shown in Fig.3, our proposed framework with ED prediction consistently outperforms CosyVoice with a higher E-MOS score, while MixedEmotion, which trains on in-domain emotional data, achieves the top score below the ground truth. In contrast, both CosyVoice and our proposed system rely solely on zero-shot emotion cloning. These results indicate that our proposed method effectively captures and replicates emotional nuances in synthesized speech, thereby enhancing emotion cloning performance.
@@
We further conducted four XAB tests, grouping 4 emotion pairs 2 : 1) ‘Anger’ vs. ‘Anxious’; 2) ‘Excited’ vs. ‘Pleasure’; 3) ‘Protected’ vs. ‘Relaxed’ ; 4) ‘Alert’ vs. ‘Surprise’. For each pair, participants were asked to correctly match the emotions, such as choosing between option A: X is Anger and Y is Anxiety and option B: X is Anxiety and Y is Anger. The results are summarized in Fig. 4. For all 4 emotion pairs, the correct match consistently exceeds 50%, with the ‘Anger vs. Anxious’ pair achieving the highest accuracy at around 83.75%. The results in Fig.4 demonstrate a positive alignment with 2 key findings in [18]:
@@
  • ‘Dominance’ is a key factor in distinguishing ‘Angry’ from ‘Anxious’, ‘Alert’ from ‘Surprised’, and ‘Relaxed’ from ‘Protected’ [^4] The first emotion in each pair involves dominance; the second emotion involves submissiveness.;
  • ‘Pleasure’ and ‘Arousal’ combine to contribute to ‘Excitement’;
@@
These observations indicate that our proposed TTS system can render emotions in accordance with psychological theories on the multidimensional nature of emotions, incorporating factors such as pleasure, arousal, and dominance. This enables our model to synthesize speech with a variety of emotional styles, enhancing both the naturalness and emotional intelligibility, and effectively bridging psychological theory with speech synthesis applications.
@@ 4 Conclusion
In this paper, we proposed an emotional TTS framework that integrates established emotion theory with language model-based speech generation. By controlling pleasure, arousal, and dominance, our proposed system synthesizes a wide range of emotional speech without relying on large-scale emotion-labeled data, while preserving naturalness in zero-shot scenarios. This work highlights the value of combining psychological insights with data-driven modeling. Future directions include exploring dynamic within-utterance emotion control and multilingual adaptation, paving the way for more versatile, affect-aware TTS solutions.

### v3

@@ 3.4 Subjective Evaluation
We conducted listening experiments with 10 native English speakers, each evaluating a total of 260 synthesized speech samples. Such participant numbers and sample sizes are standard practice in TTS research [21]. We first evaluate our proposed framework on the zero-shot emotion cloning task, where the ED is inferred from the prompt speech (‘Proposed’). We selected CosyVoice [4] as the baseline and trained it from scratch on LibriTTS to ensure a fair comparison with the proposed model. A mean opinion score (MOS) test is conducted to evaluate the speech naturalness, where participants were asked to score each sample on a 5-point scale (5: Excellent, 4: Good, 3: Satisfactory, 2: Unsatisfactory, 1: Bad). As shown in Table 2, our proposed framework achieves a score of 4.54, outperforming the baseline score of 4.36, while lower than the ground-truth score of 4.80. It shows that our proposed framework achieves better naturalness than the baseline CosyVoice.
@@
We additionally conducted a mean opinion score test for emotional intelligibility (E-MOS), where participants rated each synthesized sample based on perceived emotion. Our framework was compared against two baselines: MixedEmotion [33] and CosyVoice [4]. As shown in Fig. 3, our framework consistently achieves higher E-MOS scores than CosyVoice. MixedEmotion, which is trained on emotional data, attains the best score among systems but still falls below ground-truth speech. In contrast, both CosyVoice and our framework operate in a zero-shot setting, where no explicit emotional labels or emotional data are used during TTS training, and emotion rendering relies solely on conditioning with a reference speech prompt. Under this more challenging scenario, our method demonstrates advantages in emotional intelligibility. These results indicate that our proposed method effectively captures and replicates emotional nuances in synthesized speech, thereby enhancing emotion cloning performance.
@@
To further assess whether our system can synthesize distinguishable emotional categories, especially those with overlapping acoustic characteristics, we conducted XAB tests on four emotion pairs2: 1) ‘Anger’ vs. ‘Anxious’; 2) ‘Excited’ vs. ‘Pleasure’; 3) ‘Protected’ vs. ‘Relaxed’ ; 4) ‘Alert’ vs. ‘Surprise’. These pairs were selected because they are theoretically differentiated by specific PAD dimensions (e.g., dominance in Anger vs. Anxious), making them more challenging to disambiguate. For each pair, participants were asked to correctly match the emotions, such as choosing between option A: X is Anger and Y is Anxiety and option B: X is Anxiety and Y is Anger. The results are summarized in Fig. 4. For all 4 emotion pairs, the correct match consistently exceeds 50%, with the ‘Anger vs. Anxious’ pair achieving the highest accuracy at around 84%. The results in Fig.4 demonstrate a positive alignment with 2 key findings in [15]:
@@
  • ‘Dominance’ serves as a key factor in differentiating emotions, for example distinguishing ‘Angry’ from ‘Anxious’, ‘Alert’ from ‘Surprised’, and ‘Relaxed’ from ‘Protected’ [^4] The first emotion in each pair involves dominance; the second emotion involves submissiveness.;
  • The combination of elevated ‘Pleasure’ and ‘Arousal’ dimensions is commonly associated with ‘Excitement’;
@@
These results demonstrate that our TTS system can render emotions consistent with psychological theories of the multidimensional nature of affect, incorporating pleasure, arousal, and dominance. Unlike prior systems that rely on limited categorical labels or costly dimensional annotations and thus struggle to capture the full range of emotions, our model leverages continuous dimensions to synthesize speech with diverse emotional styles. This enhances both naturalness and emotional intelligibility, effectively bridging psychological theory with practical speech synthesis.
Fig. 4: XAB test results for four synthesized emotion pairs, used to evaluate the intelligibility of perceived emotions and the effectiveness of emotion control in the proposed system.
@@ 4 Conclusion
In this paper, we proposed an emotional TTS framework that integrates established emotion theory with language model-based speech generation. By controlling pleasure, arousal, and dominance, our proposed system synthesizes a wide range of emotional speech without relying on large-scale emotion-labeled data, while preserving naturalness in zero-shot emotion cloning scenarios. This work highlights the value of combining psychological insights with data-driven modeling. Future directions include exploring dynamic within-utterance emotion control and multilingual adaptation, paving the way for more versatile, affect-aware TTS solutions.